### v1

# Learning Disentangled Stain and Structural Representations for Semi-Supervised Histopathology Segmentation

Abstract
Accurate gland segmentation in histopathology images is essential for cancer diagnosis and prognosis. However, significant variability in Hematoxylin and Eosin (H&E) staining and tissue morphology, combined with limited annotated data, poses major challenges for automated segmentation. To address this, we propose Color-Structure Dual-Student (CSDS), a novel semi-supervised segmentation framework designed to learn disentangled representations of stain appearance and tissue structure. CSDS comprises two specialized student networks: one trained on stain-augmented inputs to model chromatic variation, and the other on structure-augmented inputs to capture morphological cues. A shared teacher network, updated via Exponential Moving Average (EMA), supervises both students through pseudo-labels. To further improve label reliability, we introduce stain-aware and structure-aware uncertainty estimation modules that adaptively modulate the contribution of each student during training. Experiments on the GlaS and CRAG datasets show that CSDS achieves state-of-the-art performance in low-label settings, with Dice score improvements of up to 1.2% on GlaS and 0.7% on CRAG at 5% labeled data, and 0.7% and 1.4% at 10%. Our code and pre-trained models are available at https://github.com/hieuphamha19/CSDS.
Keywords: Histopathology Imaging, Semi-Supervised Learning, Semantic Segmentation.
1 Introduction
Histopathological image analysis is of paramount importance in cancer diagnosis and prognosis, especially in tasks such as gland segmentation for colorectal cancer. Images are usually stained with Hematoxylin and Eosin (H&E). This staining can create significant color differences due to variations in staining methods, slide preparation, and scanning equipment used. The differences in color and structure of tissue images create considerable obstacles for automated segmentation systems, especially when there is a lack of labeled data. In medical imaging, however, obtaining such labels is expensive, time-consuming, and subject to inter-observer variability.
Semi-supervised learning (SSL) offers an effective strategy to improve segmentation performance by leveraging a small amount of labeled data alongside a larger pool of unlabeled samples. However, existing SSL methods often overlook the unique visual complexity of histopathological images, which arises from two entangled yet semantically distinct sources: staining variations and tissue morphology. In particular, Hematoxylin and Eosin (H&E) staining introduces substantial inter-sample color variation due to inconsistencies in staining protocols and scanning devices, while structural variations reflect biologically meaningful changes in tissue architecture, including malignant progression. Prior works in semi-supervised segmentation, such as pseudo-labeling Zhang et al. (2021), consistency regularization Sohn et al. (2020), and teacher-student frameworks Tarvainen and Valpola (2017); Yu et al. (2019); Pham et al. (2025), generally process the image holistically and fail to disentangle stain and structure cues. This lack of targeted representation learning can limit their effectiveness in histopathology, where isolating stain-specific and structure-specific information is crucial for robust and interpretable segmentation.
To overcome the limitations of existing semi-supervised segmentation approaches, we propose in this work a novel framework termed Color-Structure Dual-Student (CSDS). The key innovation of CSDS lies in its explicit decoupling of color and structural information through the use of two specialized student networks. One student is trained with color-augmented images to effectively model chromatic variations, while the other is trained with geometrically transformed images to emphasize structural cues. A shared teacher model, updated using an Exponential Moving Average (EMA) of the student weights, then serves as a pseudo-label generator and provides supervision in an uncertainty-aware manner.
To further enhance pseudo-label reliability, we introduce color-aware and structure-aware uncertainty estimation strategies. These mechanisms leverage prediction entropy and domain-specific priors to adaptively weight the contributions of each student during training, thereby improving the quality and robustness of the supervision signal. We validate the effectiveness of CSDS through comprehensive experiments conducted on two histopathological image segmentation benchmarks, GlaS Sirinukunwattana et al. (2017) and CRAG Graham et al. (2019), under low-label regimes. The results demonstrate that CSDS consistently outperforms existing state-of-the-art methods, establishing a new benchmark for semi-supervised medical image segmentation.
2 Methodology
2.1 Preliminary
Problem Settings. Following standard semi-supervised segmentation settings, we denote the labeled set as $\mathcal{D}_{l}=\{(x_{i}^{l},y_{i}^{l})\}_{i=1}^{N}$, and a larger unlabeled set $\mathcal{D}_{u}=\{(x_{k}^{u})\}_{m=1}^{M}$ ($N$: labeled samples, $M$: unlabeled samples, $M\gg N$). Here, $x_{i}^{l}$ and $x_{k}^{u}$ are input images from the labeled and unlabeled sets, respectively, and $y_{i}^{l}$ is the corresponding one-hot groundtruth for labeled sample $x_{i}^{l}$.
Overall Pipeline. We propose a Color-Structure Dual-Student (CSDS) framework, an extension of the Mean Teacher paradigm Tarvainen and Valpola (2017). Our framework is built upon the intuition that color and structure represent two complementary sources of variation in histological data: while color distributions often shift across staining protocols, structural patterns exhibit large inter-sample variability due to tissue deformation and cancer heterogeneity. CSDS comprises two student networks with shared architecture: the color student, trained on inputs augmented by histogram matching and color jittering to capture both in-distribution and out-of-distribution color shifts; and the structure student, optimized on elastically deformed images to model structural variability. A shared teacher network is updated via Exponential Moving Average (EMA) of the mean weights from both students. In Section 2.2 we introduce two complementary uncertainty estimation modules: Color uncertainty and structure uncertainty, which quantify prediction confidence via entropy in color-ambiguous and structurally-uncertain regions, respectively. These uncertainty maps are subsequently used to modulate the teacher’s supervision signals during the pseudo-labeling process. The overall pipeline of our method is illustrated in Figure 1.
Figure 1: Overview of our Color-Structure Dual-Student (CSDS) framework with dual students for color and structure, and a shared EMA-updated teacher. Output uncertainty maps are used to adaptively weight pseudo-labels during training.
2.2 Uncertainty Estimation for Color-Structure Uncertainty Module
In this section, we present the detailed design of the Color Uncertainty and Structure Uncertainty modules. Uncertainty map is initially computed from the teacher’s predictions using Shannon entropy Shannon (1948) and is further refined by incorporating visual features extracted from input images, guided by color and structural indicators Zou et al. (2023).
2.2.1 Entropy-Based Uncertainty Estimation
Given the teacher’s output logits $\mathbf{z}\in\mathbb{R}^{C\times H\times W}$, we compute per-pixel softmax probabilities:
$\mathbf{p}(x)=\operatorname{softmax}(\mathbf{z}(x)),$ (1)
where $z(x)$ denotes the logit at pixel $x$. Uncertainty is then estimated as:
$U(x)=-\sum_{c=1}^{C}p_{c}(x)\log(p_{c}(x)+\varepsilon),$ (2)
where $\varepsilon$ is a small constant added for numerical stability. This produces a dense uncertainty map that reflects the teacher’s predictive confidence.
2.2.2 Color-Aware Uncertainty Modulation
Within the domain of pathology imaging, local color diversity often reflects meaningful semantic variation Xia et al. (2024). Given the input RGB image $I\in\mathbb{R}^{3\times H\times W}$, we estimate chromatic complexity by computing the per-pixel inter-channel variance:
$\sigma^{2}(x)=\operatorname{Var}_{c\in\{R,G,B\}}[I_{c}(x)].$ (3)
To enforce spatial coherence, the variance map is smoothed using a Gaussian filter $\mathcal{G}_{3\times 3}$, or approximated via average pooling for computational efficiency:
$V(x)=\mathcal{G}_{3\times 3}(\sigma^{2}(x)),\quad\hat{V}(x)=\frac{V(x)}{\max_{x\in\Omega}V(x)+\varepsilon}.$ (4)
A binary mask highlights regions with high color diversity:
$M_{\mathcal{C}}(x)=\mathbb{I}[\hat{V}(x)>\tau_{\text{color}}].$ (5)
We then modulate the base uncertainty map $U(x)$ by amplifying values in chromatically distinctive regions:
$\widetilde{U}_{\mathcal{C}}(x)=U(x)\cdot\left(1+\lambda_{\mathcal{C}}\cdot M_{\mathcal{C}}(x)\right),$ (6)
where $\lambda_{\mathcal{C}}\in[0,1]$ controls the degree of modulation. This adjustment selectively increases uncertainty in visually rich areas where semantic ambiguity may be elevated.
2.2.3 Structure-Aware Uncertainty Modulation
Uncertainty tends to rise near object boundaries and regions of strong visual transition, where label ambiguity is more likely Yang et al. (2025). To model this, we compute approximate spatial gradients using finite differences:
$G_{h}(x)=\left|I(x+1,y)-I(x,y)\right|,\quad G_{v}(x)=\left|I(x,y+1)-I(x,y)\right|.$ (7)
An edge magnitude map is formed and normalized:
$E(x)=\frac{1}{3}\left(G_{h}(x)+G_{v}(x)\right),\quad\hat{E}(x)=\frac{E(x)}{\max_{x\in\Omega}E(x)+\varepsilon}.$ (8)
We then identify structurally salient regions:
$M_{\mathcal{S}}(x)=\mathbb{I}[\hat{E}(x)>\tau_{\text{structure}}].$ (9)
Finally, the uncertainty map is further refined by emphasizing these structure-rich areas:
$\widetilde{U}_{\mathcal{S}}(x)=\widetilde{U}_{\mathcal{C}}(x)\cdot\left(1+\lambda_{\mathcal{S}}\cdot M_{\mathcal{S}}(x)\right),$ (10)
where $\lambda_{\mathcal{S}}\in[0,1]$ controls the strength of modulation. The final uncertainty map $\widetilde{U}_{\mathcal{S}}$ integrates both chromatic and structural information, yielding a perception-aware enhancement that emphasizes regions of higher epistemic risk.
2.3 Training Procedure
We denote the teacher network as $\mathcal{T}$, the color student as ${\mathcal{C}}$, and the structure student as ${\mathcal{S}}$. In the supervised phase, the loss is computed as:
$\mathcal{L}_{\text{sup}}=\mathcal{L}_{\text{CE+Dice}}({\mathcal{C}}(x^{l}),y^{l})+\mathcal{L}_{\text{CE+Dice}}({\mathcal{S}}(x^{l}),y^{l}),$ (11)
[…]
$\mathcal{L}_{\text{unsup}}=$ $\;\widetilde{U}_{\mathcal{C}}(x^{u})\cdot\mathcal{L}_{\text{CE+Dice}}({\mathcal{C}}(\mathcal{A}_{\mathcal{C}}(x^{u})),\mathcal{T}(x^{u}))$ (12)
$+\widetilde{U}_{\mathcal{S}}(x^{u})\cdot\mathcal{L}_{\text{CE+Dice}}({\mathcal{S}}(\mathcal{A}_{\mathcal{S}}(x^{u})),\mathcal{T}(\mathcal{A}_{\mathcal{S}}(x^{u})))$
Finally, the total training loss combines both objectives:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{sup}}+\lambda_{\text{unsup}}\mathcal{L}_{\text{unsup}},$ (13)
where $\lambda_{\text{unsup}}$ is a weighting factor that controls the unsupervised loss contribution.
3 Experiments
3.1 Datasets
We evaluated our proposed method on two publicly available histopathological benchmarks. GlaS Sirinukunwattana et al. (2017) is a gland segmentation dataset for colorectal adenocarcinoma across various cancer stages, containing 165 images (37 benign, 48 malignant, and 80 test) with resolutions around $775\times 522$ or $589\times 453$. CRAG Graham et al. (2019) targets gland segmentation in colon histopathology, comprising 213 annotated images, mostly sized $1512\times 1516$. It is particularly challenging due to substantial variations in gland morphology and staining. Following prior works, 20% of the images were used as a fixed test set, and the remaining 80% were split into five folds for cross-validation on both datasets.
3.2 Implementation Details
All experiments were conducted on a single NVIDIA RTX 3060 GPU with Python 3.11, PyTorch 2.5, and CUDA 12.2. For both GlaS and CRAG datasets, we adopted DeepLabV3+ with a ResNet-101 backbone as the segmentation model. Models were trained for 80 epochs on GlaS and 120 epochs on CRAG until convergence, with all images resized to $256\times 256$.
To enhance generalization, we applied shared augmentations across branches, including random horizontal/vertical flips and rotations. In addition, each student received branch-specific augmentations. The Color Student was trained with color-focused transformations using a OneOf strategy: either random color jittering (with brightness, contrast, saturation set to 0.3 and hue to 0.05) or histogram matching Ferrero et al. (2024), which preserves structural integrity while increasing color robustness. In contrast, the Structure Student was augmented with elastic deformation Faryna et al. (2021), perturbing spatial structures while retaining color, thereby enhancing sensitivity to boundaries and textures.
We used AdamW with a learning rate of $1\times 10^{-4}$, $\beta=(0.9,0.999)$, $\epsilon=1\times 10^{-8}$, and a weight decay of 0.05. A batch size of 4 empirically yielded the best performance. During inference, the student model with the best validation performance is selected for generating predictions.
3.3 Comparison with State-of-the-arts
We benchmark the proposed method against eight SOTA semi-supervised segmentation methods, covering both general-purpose and histopathology-specific approaches Tarvainen and Valpola (2017); Yu et al. (2019); Sohn et al. (2020); Chen et al. (2021); Luo et al. (2022); Wang et al. (2023); Sun et al. (2024); Nguyen et al. (2025). We report the average performance across five folds along with standard deviation.
Labeled Ratio Method GlaS CRAG
Dice $\uparrow$ Jaccard $\uparrow$ Dice $\uparrow$ Jaccard $\uparrow$
100% Fully-Supervised $\mathbf{90.84\pm 0.28}$ $\mathbf{83.72\pm 0.42}$ $\mathbf{86.36\pm 0.73}$ $\mathbf{77.38\pm 1.05}$
5% Supervised Baseline $72.31\pm 3.19$ $58.82\pm 3.43$ $63.87\pm 3.29$ $50.73\pm 2.78$
Mean Teacher (NeurIPS 2017) $77.23\pm 3.08$ $64.59\pm 3.34$ $71.71\pm 2.06$ $58.78\pm 2.35$
UAMT (MICCAI 2019) $79.43\pm 3.28$ $67.85\pm 3.75$ $72.01\pm 2.45$ $59.53\pm 2.30$
Fixmatch (NeurIPS 2020) $79.15\pm 1.75$ $67.66\pm 2.21$ $71.67\pm 3.36$ $58.83\pm 3.56$
CPS (CVPR 2021) $79.55\pm 2.01$ $67.84\pm 3.21$ $74.34\pm 2.19$ $61.36\pm 2.58$
CT (MIDL 2022) $79.86\pm 1.49$ $68.13\pm 3.64$ $71.84\pm 3.93$ $58.34\pm 4.79$
CCVC (CVPR 2023) $80.84\pm 1.75$ $68.88\pm 2.28$ $73.28\pm 1.87$ $60.48\pm 2.17$
CorrMatch (CVPR 2024) $79.85\pm 1.96$ $67.79\pm 2.79$ $69.08\pm 2.31$ $55.39\pm 2.88$
FDCL (AAAI-25 Bridge Program) $81.64\pm 1.08$ $70.15\pm 1.52$ $74.55\pm 1.51$ $61.93\pm 1.84$
CSDS (Ours) $\mathbf{82.86\pm 1.24}$ $\mathbf{72.01\pm 1.81}$ $\mathbf{75,25\pm 1,54}$ $\mathbf{63,00\pm 1,67}$
10% Supervised Baseline $75.64\pm 3.99$ $62.48\pm 4.11$ $71.80\pm 3.11$ $58.91\pm 3.20$
Mean Teacher (NeurIPS 2017) $79.68\pm 3.33$ $68.08\pm 4.11$ $75.36\pm 2.17$ $62.77\pm 2.61$
UAMT (MICCAI 2019) $84.05\pm 0.78$ $74.20\pm 1.65$ $75.48\pm 3.16$ $63.05\pm 3.42$
Fixmatch (NeurIPS 2020) $81.13\pm 2.05$ $69.40\pm 3.71$ $74.87\pm 3.20$ $62.30\pm 3.38$
CPS (CVPR 2021) $84.29\pm 0.44$ $73.89\pm 0.88$ $78.08\pm 1.32$ $66.11\pm 1.79$
CT (MIDL 2022) $82.67\pm 1.15$ $71.65\pm 2.98$ $73.68\pm 1.86$ $60.44\pm 2.19$
CCVC (CVPR 2023) $83.78\pm 2.31$ $73.52\pm 2.99$ $74.97\pm 1.40$ $62.30\pm 1.76$
CorrMatch (CVPR 2024) $83.27\pm 2.23$ $72.59\pm 2.01$ $74.90\pm 1.41$ $61.93\pm 2.35$
FDCL (AAAI-25 Bridge Program) $84.35\pm 1.12$ $74.45\pm 1.14$ $76.28\pm 2.29$ $63.92\pm 2.88$
CSDS (Ours) $\mathbf{85.06\pm 0.57}$ $\mathbf{74.87\pm 0.82}$ $\mathbf{79.50\pm 0.88}$ $\mathbf{68.14\pm 1.18}$
Table 1: Quantitative results on GlaS-2017 and CRAG-2019 datasets on two labeled ratios. We highlight the best performance for each metric in bold.
3.4 Quantitative Results
As shown in Table 1, CSDS consistently outperforms previous SOTA methods under limited supervision on both GlaS and CRAG datasets. On the Glas dataset, with only 10% labeled data, CSDS achieves a Dice score of $82.86\pm 1.24$ and a Jaccard index of $72.01\pm 1.81$, which surpasses strong co-training baselines like FDCL and CPS. In 5% labeled ratio, CSDS narrows the Dice gap to full supervision to just 0.81%, while significantly outperforming teacher-student methods (e.g., Mean Teacher) and self-training methods (e.g., FixMatch). CSDS generalizes well across datasets, as evidenced by consistently outperforming all compared methods under all settings on CRAG, which presents more irregular and challenging nuclei structures.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Image GT UAMT CPS CCVC CorrMatch FDCL Ours
Figure 2: Qualitative results for different semi-supervised methods on 10% labeled data from two datasets. Rows 1 to 3 correspond to GlaS, while Rows 4 to 5 correspond to CRAG. The red boxes emphasize the differences in the results.
3.5 Qualitative Results
Figure 2 shows qualitative results on GlaS (Rows 1–3) with 10% labeled data. While most methods capture overall morphology, earlier approaches like UAMT and CPS struggle with boundary precision and separating clustered glands. FDCL and CorrMatch improve but still show artifacts. In contrast, CSDS produces cleaner, more accurate masks with sharper boundaries, aligning with the quantitative results in Table 1. Rows 4 and 5 show results on CRAG. While most methods fail to capture complete nuclei contours, causing under-segmentation or fragmented boundaries, CSDS achieves the most consistent results with clearer boundaries and better shape preservation.
3.6 Ablation study
Effect of Model Components. As shown in Table 2, removing either the Color or Structure Student leads to clear performance drops. Using both together significantly improves results, confirming their complementary roles in enhancing representation learning and segmentation.
SupRatio Teacher Color Student Structure Student Dice (%) Jaccard (%)
5% $81.10\pm 2.32$ $69.53\pm 2.97$
$80.35\pm 3.47$ $68.75\pm 4.11$
$\mathbf{82.86\pm 1.24}$ $\mathbf{72.01\pm 1.81}$
10% $84.37\pm 1.27$ $74.02\pm 1.70$
$84.07\pm 1.55$ $73.75\pm 2.03$
$\mathbf{85.06\pm 0.57}$ $\mathbf{74.87\pm 0.82}$
Table 2: Ablation study using different Teacher, Color, Structure Student combinations.
Effect of EMA Strategies. As shown in Table 3, the Mean strategy, which updates the teacher model by averaging the parameters of both students, consistently achieves the best segmentation performance. In contrast, Alternate, which updates the teacher by switching between the two branches each epoch, and Best student only, which updates using the student with higher validation performance, perform worse. These results suggest that aggregating knowledge from both students yields more stable and informative teacher updates by reducing potential variance and bias introduced by a single student branch.
SupRatio EMA Strategy Dice (%) Jaccard (%)
0.05 Alternate $81.80\pm 2.38$ $70.44\pm 3.19$
Best student only $81.60\pm 1.88$ $70.13\pm 2.56$
Mean $\mathbf{82.86\pm 1.24}$ $\mathbf{72.01\pm 1.81}$
0.10 Alternate $84.56\pm 1.40$ $74.30\pm 1.93$
Best student only $84.34\pm 1.64$ $74.03\pm 2.28$
Mean $\mathbf{85.06\pm 0.57}$ $\mathbf{74.87\pm 0.82}$
Table 3: Ablation study of different EMA strategies under varying supervision ratios.
4 Conclusion
In this work, we introduced Color-Structure Dual-Student (CSDS), a novel semi-supervised segmentation framework tailored for histopathology images. By explicitly decoupling color and structural information into two specialized student networks, guided by a shared teacher model, CSDS effectively captures domain-specific variations. Our uncertainty-aware training strategy further improves pseudo-label reliability through color-adaptive and structure-adaptive weighting. Extensive experiments on GlaS and CRAG datasets under limited-label regimes demonstrate that CSDS outperforms strong baselines and recent state-of-the-art methods. These findings underscore the effectiveness of explicitly modeling visual cues unique to histopathology in semi-supervised segmentation.